OpenStudy (anonymous):

Find the slope-intercept form of the line that passes through given the point and has the given slope. (1, 3) m = 2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

so yeah y-y1=m*(x-x1) m=slope and (x1, y1) are given coordinates

OpenStudy (alexandervonhumboldt2):

oh wait thaT ANOTHER FORM

OpenStudy (alexandervonhumboldt2):

y=ax+b where a is the given slope then substitue oints into the equation and solve for b

OpenStudy (anonymous):

Slope-intercept (1, 3) m = 3